An amoeba has 3.3 x10^16 protons and a net charge of 1.1 pC. How many fewer electrons are there than protons?If you paired them up, what fraction of the protons would have no electrons?
Answer #1

A)

N - n = (1.1 x 10^-12)/(1.6 x 10^-19) = 6.88 x 10^6

B)

Fraction = (6.88 x 10^6)/(3.3 x 10^16) = 2.08 x 10^-10

Comment in case any doubt please rate my answer....